INTENSITIES, SELF- AND AIR-BROADENING, AND PRESSURE-INDUCED LINE SHIFTS IN THE 2ν_{2} BAND OF HCN

V. MALATHY DEVI, \underline{D. CHRIS BENNER}, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795; M. A. H. SMITH, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; S. W. SHARPE and R. L. SAMS, Mail Stop K8-88, Battelle Blvd., Pacific Northwest National Laboratory, P. O. Box 999, Richland, WA 99352.

High-resolution infrared absorption spectra of HCN in the 2ν_{2} band region near 1411 cm^{−1} have been recorded with the Bruker IFS 120 HR Fourier transform spectrometer located at Pacific Northwest National Laboratory. The measurements were all made at ambient temperatures between 24.7 and 24.9^{∘}C, although the variation during the course of recording each spectrum was less than ±0.01^{∘}C. Two stainless steel absorption cells that were electroplated and gold coated to minimize reactivity were used to obtain the spectra; their path lengths were 0.958 and 19.95 cm. Four spectra of highpurity (99.8%) HCN together with three spectra of lean mixtures (∼ 3%) of HCN in dry air were simultaneously fit using a multispectrum nonlinear least squares procedure.^{a} The analysis yielded room-temperature values for absolute intensities, self- and air-broadening coefficients and self- and air-induced pressure shift coefficients for numerous lines in the 2ν_{2} band of H^{12}C^{14}N. Since there are no previous measurements of broadening and shift parameters in the 2ν_{2} band, our results will be compared with values recently determined in the ν_{1} band of H^{12}C^{14}N^{b} and with current HITRAN values.^{c}

aD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).

bC. P. Rinsland et al., JQSRT, in press (2003).

cL. S. Rothman et al., JQSRT, in press (2003).